OpenStudy (anonymous):

Solve the equation. log(x+7) - logx=3

ganeshie8 (ganeshie8):

Hint : \[\large \log( a) - \log (b) = \log \left(\frac{a}{b}\right)\]

OpenStudy (mimi_x3):

The first step like Ganeshie has stated is to combine the 2 logarithmic functions. Can you attempt to use Ganeshie's hint?

OpenStudy (anonymous):

no, i dont understand were i need to plug in the numbers

ganeshie8 (ganeshie8):

what do you know about logs ?

OpenStudy (anonymous):

not much really

ganeshie8 (ganeshie8):

thats good because we can start fresh :)

ganeshie8 (ganeshie8):

`logs` are same as `exponents` you might be knowing that \(2\times 2\times 2\) equals \(8\) and we write it in short form as : \[\large 2^3 = 8\]

ganeshie8 (ganeshie8):

another way to write the same thing is : \[\large \log_2 8 = 3\]

OpenStudy (anonymous):

oh okay

ganeshie8 (ganeshie8):

let me ask you a question

ganeshie8 (ganeshie8):

write below in log form : \[\large 3^4 = 81\]

OpenStudy (anonymous):

\[\log _{3}81= 4\]

OpenStudy (anonymous):

no?

ganeshie8 (ganeshie8):

Perfect!

ganeshie8 (ganeshie8):

Lets look at our original problem. Applying the first property, \[\log(x+7) - \log x=3\] becomes \[\log\left(\frac{x+7}{x}\right)=3\]

ganeshie8 (ganeshie8):

Btw, the default base is \(10\) : \[\log_{10}\left(\frac{x+7}{x}\right)=3\] changing this to exponent form you get : \[\frac{x+7}{x} = 10^3\]

ganeshie8 (ganeshie8):

\[\frac{x+7}{x} = 1000\]

ganeshie8 (ganeshie8):

\[x+7 = 1000x\] isolate \(x\)